OpenStudy (anonymous):

factor out the trinomial. h=-2tsquared+87t+90

OpenStudy (anonymous):

well i know from there what you do is \[-(2t ^{2}-87t-90)\] then you multiply 90 and 2t squared which gives you -180 then what you have to do is think what times what vies me 180 but also when u add/subtract them it gives you 87 and i dont know those two number

OpenStudy (anonymous):

Just use the quadratic equation \[t = [-b \pm \sqrt{b^2-4ac}]/2a\] Where a=-2, b=87, c=90. That will give you your two values for t when h=0 allowing you to rewrite the equation as \[h = (t-t_1)(t-t_2)\]

OpenStudy (anonymous):

i never learned that i just need two number that multiply and equal 180 but when u add/subtract those two numbers in equals 16

OpenStudy (anonymous):

Are you sure the middle term isn't 88t?

OpenStudy (anonymous):

im sure

OpenStudy (anonymous):

Then you don't get nice numbers. If it were 88t you'd have: h=-(2t+2)(t-45) As it is here, you get h=(t + 1.0109863582402774)(t-44.510986358240274)

OpenStudy (anonymous):

oh woah i will just ask my teacher tomorrow

OpenStudy (anonymous):

thank you though!

OpenStudy (anonymous):

oh and also what times what equals 1536 but when u add/subtract them it eguals -16